OpenStudy (anonymous):

Question 1 The table represents data collected on the time spent studying (in minutes) and the resulting test grade. Time Spent Studying (min) 52 37 31 9 26 40 22 10 45 34 19 60 Grade on Test 95 84 72 58 77 86 72 43 90 81 62 98 Find the line of best fit for the data either by hand or using technology. Explain your method. Find the predicted score for each time listed in the table.

OpenStudy (kropot72):

How do you intend to do this, by hand or with the use of a calculator?

OpenStudy (anonymous):

hand or calc whichever is easier lol

OpenStudy (kropot72):

Do you have a calculator that can do regression?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Then you need to input the data, having selected the type of statistical calculation. On my calculator I selected 'Paired variable (X, Y), linear regression (y = A + Bx)'.

OpenStudy (anonymous):

are those the steps?

OpenStudy (kropot72):

Yes. Your link gives the solution to the question that you posted, assuming linear regression. What help do you need?

OpenStudy (anonymous):

so by finding the best fit for the data it is asking me to the find the regression line? if so, do i write all those steps the link gave to me and that will be my answer?

OpenStudy (kropot72):

The equation for the regression line is: y = 45.547 + 0.965x where y is the predicted grade and x is the time spent studying (in minutes). For example the predicted score for a study time of 40 minutes is found from: y = 45.547 + (0.965 * 40) = 84.147 It is not clear whether or not you need to state all the details of the calculation done by the calculator. However the question requires you to calculate the predicted grade for each time given in the table.

OpenStudy (anonymous):

thankyou :)

OpenStudy (kropot72):

You're welcome :)